# Training LLMs over Neurally Compressed Text

### v1

@@ 5.2 Transformers struggle to learn Arithmetic Coding
Task Accuracy Cross Entropy
Decompression $76.98$ $0.751\pm 0.005$
@@
Compression $1.7$ $2.489$
To directly diagnose the ability to track Arithmetic Coding, we format AC compression and decompression as sequence-to-sequence tasks. The input provides the model with the true text, so we expect a model that is able to learn Arithmetic Coding should achieve an accuracy of $100$. We compress sequences of $1{,}024$ bytes using M1 and Arithmetic Coding. [^13] We use shorter raw text sequences to keep the final sequence length of inputs + targets manageable. We concatenate the bytes and AC output tokens to create the compression task. For the decompression task, we simply flip the order—AC output tokens first and then bytes. The target tokens (bytes or tokens) are shifted by the input vocabulary size, ensuring that they have distinct values. We use a decoder-only transformer as our model with a causal attention mask, i.e., even during the input sequence, future tokens are hidden from the model. We train models with $113$m parameters. Loss, gradients, and evaluation metrics are only computed on the target tokens.
In the decompression task, the target tokens are bytes. By ignoring the inputs and just modeling the outputs, the decompression model can achieve decent performance without actually leveraging the input data. To control for this, we also train a byte-level language model baseline on the same sequence-to-sequence data, excluding the input tokens. If the decompression model is actually learning to decompress Arithmetic Coding, we would expect stronger performance than the byte-level baseline. As we see in Table 5, the baseline model, which does not see the input tokens, has the same performance as the decompression model. [^14] The slight gain is statistically insignificant, $(p=0.07)$. Clearly, the models trained for decompression are not actually learning to do decompression.
@@ 6.1 EqualInfoAC is less stable and less semantic than SentencePiece
First, we observe that SentencePiece produces a relatively stable text $\rightarrow$ token mapping. [^17] See Appendix L for some corner cases where this is not the case. For example, “elephants” appears three times in the sentence, and maps stably to the same two-token sequence in all cases: [ elephant][s]. Similarly, both occurrences of “African” map to the same token: [ African]. By comparison, the EqualInfoAC tokenization is relatively unstable, with each occurrence of these words being segmented in a different way and yielding a different token sequence.
@@
Despite these differences, there is an important similarity between SentencePiece and EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$: they are both stable in the token $\rightarrow$ text direction. That is, a given token ID, e.g., token #$500$, will always map to the same output text. This “transparent decoding” property likely makes it easier for a downstream model to learn over these tokens. [^18] Padding to reach a specific window size can require extra computation to discern between padding and characters that compress to all zeros, however we find in Appendix I that it is not an issue for M2 models.
@@
When we move to versions of EqualInfoAC that contain multiple tokens per window, such as EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}256]$, this transparency is destroyed for all non-initial tokens within a window. This is illustrated in Table 8. When the same token appears window-initially in different contexts, we see the window text has a stable prefix—e.g., token #$151$ always maps to the prefix “le-”. However, when occurring as the second token within a two-token window, there are no apparent correspondences between window text. [^19] A repeated text substring that happens to be aligned with a window multiple times is one of the few cases where the second token will represent the same text. As EqualInfoAC window length increases, the proportion of tokens that are stable decreases. This may explain the observed difficulty of learning over longer windows. The window text for all instances of these tokens can be seen in Appendix M.
@@
Note that Table 8 examines window $\rightarrow$ text, as opposed to token $\rightarrow$ text correspondences. This is because for multi-token windows, the mapping from tokens to text is not well defined. More specifically, each character maps to a particular subsequence of the compressed bitstream, but these may not align with token boundaries. [^20] This can be a source of instability, even in window-initial tokens, see Appendix L. Fig. 7 illustrates the mapping between characters, bits, and tokens. We find that many windows contain a character (shown in purple) whose bits are split across two $8$-bit tokens.
Fig. 7 also highlights that window-initial characters are not being well compressed, with the window-initial token often only covering one or two characters. This is due to our EqualInfoAC procedure fully resetting M1’s context at every window boundary. With no context, M1 cannot make confident predictions, leading to more bits being needed to represent the initial character. On the positive side, this setup guarantees that a window can be decoded in isolation, which should aid learning. However it is worth exploring in future work whether maintaining some M1 context across windows could improve the compression ratio without hurting learnability.
@@ 6.2 AC decoding is learned step-by-step
The second trend concerns the accuracy reached at each position. Here, we observe an increase in accuracy from #1 < #2 < #3, followed by a decrease from #3 < #4 < #5 and so on. [^22] The final token #8 also fits this trend when looking at the increase over non-trivial accuracy. The raw accuracy in this position is higher than previous tokens #4–7, due to the skewed distribution introduced by window-final padding. We interpret the increase across the first three positions as due to the benefit of extra leftward context. This is akin to the initial byte in a word being harder to predict than the following bytes. The decreasing performance at tokens #4 and beyond suggests the model is unable to track AC decompression indefinitely. While the model clearly learns to decompress longer sequences as training progresses, reliably decoding past $32$ bits of AC output appears to be a challenge.
@@ 6.3 Learnable distributions are less uniform
A well-known result in the compression literature is that there can be no recursive compression [45]. The compression algorithm removes information captured by its model, resulting in a uniform output that appears random to the original model. However, our setting is not recursive compression. Instead, a separate and larger model is trained on the compressed output, which should be able to capture new patterns in the bitstream.
Despite this, the output of compression using M1 appears very uniform, as evidenced by the minimal gains from modeling the unigram token distribution in Table 3. Therefore, it seems reasonable that this uniformity could make it hard for M2 to learn (as all patterns must be contextual). We investigate this by plotting the KL divergence [39] between the observed empirical distribution and a uniform distribution for different segmentations of the bitstream. If the underlying distribution of bits was truly random and independent, then the distribution of unigrams for some bitstream segmentation should remain uniform as $p(b_{i},\dots,b_{i+n})=\prod_{j=i}^{i+n}(p(b_{j}))$ and therefore the KL divergence should remain close to zero. On the other hand, if the distribution diverges from uniform, there is contextual information to be learned when training an LLM to model $p(b_{n}|b_{i},\dots,b_{i+n-1})$.
@@
As a baseline, we used the cryptographic secrets package in Python to generate bitstreams that should be truly random and independent. As such, the KL divergence should remain at 00 when segmented in the same way as the compressed data. The reason this does not hold in Fig. 9 is that the maximum likelihood estimate of entropy, $\hat{H}=-\sum_{x\in\hat{\mathcal{X}}}\hat{p}(x)\log_{2}\hat{p}(x)$, is negatively biased [48]. In Fig. 13 we see that when using a Miller-Madow estimator [46] to correct for this bias, the expected KL of 00 is well within sampling noise bounds. To account for noise in the entropy estimation, we plot $90$th percentile intervals of the KL divergence between the observed entropy from $100$ disjoint samples of the data and the uniform distribution. [^23] As the number of bits in a segmentation grow, the vocabulary size increases exponentially, requiring many more samples. Thus we expect noise in the entropy estimate to grow with $n$. This holds, but it is obfuscated by the log scaling in Fig. 9. In fact, the magnitude of the noise for settings such as GZip and EqualInfoAC is larger than for AC or RNG. This noise behavior is seen in Fig. 12. See Appendix J for more information on entropy estimation and bias correction.
@@ 7 Conclusion
We have shown there is promise in the idea of training LLMs over neural-compressed text. In the best case, this will allow training over text that is better compressed than standard subword token sequences, while maintaining learnability. This an appealing prospect, as models that read and write more text per token are more efficient to train and serve, and can model longer dependencies.
@@
While bespoke compression methods have developed around different modalities (e.g., text, audio, images, video) and different applications (e.g., delta-of-delta for regular repeating timestamps [50]), to our knowledge, no efficient compression methods have been designed specifically for use as LLM tokenizers. We are optimistic that future work will create such methods. Compared to today’s subword tokenizers, we expect these methods (i) will deliver higher compression rates, (ii) will come closer to equal information per token, thus allocating compute more effectively, and (iii) will give models a more direct view of the underlying raw text, thus helping on spelling and pronunciation tasks. As a tradeoff, we expect these neural tokenizers will be somewhat less stable in their text $\leftrightarrow$ token mapping, but perhaps not so unstable as our approach here. In particular, we think it is worth exploring methods under which a given word typically maps to a relatively small number (tens not thousands) of relatable token sequences.

### v2

@@ 5.2 Transformers struggle to learn Arithmetic Coding
Compression Task Accuracy Cross Entropy
AC$[v\mathord{=}256]$ Decompression $76.98$ $0.751\pm 0.005$
@@
EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}256]$ Decompression $95.30$ $0.141\pm 0.002$
Byte Level LM $76.83$ $0.756\pm 0.001$
AC$[v\mathord{=}256]$ Compression $1.7$ $2.489$
To directly diagnose the ability to track Arithmetic Coding, we format AC compression and decompression as sequence-to-sequence tasks. The input provides the model with the true text, so we expect a model that is able to learn Arithmetic Coding should achieve an accuracy of $100$. We compress sequences of $1{,}024$ bytes using M1 and Arithmetic Coding. [^13] We use shorter raw text sequences to keep the final sequence length of inputs + targets manageable. We concatenate the bytes and AC output tokens to create the compression task. For the decompression task, we simply flip the order—AC output tokens first and then bytes. The target token IDs (bytes or tokens) are shifted by the input vocabulary size, ensuring that they have distinct values. We use a decoder-only transformer as our model with a causal attention mask, i.e., even during the input sequence, future tokens are hidden from the model. We train models with $113$m parameters. Loss, gradients, and evaluation metrics are only computed on the target tokens.
In the decompression task, the target tokens are bytes. By ignoring the inputs and just modeling the outputs, the decompression model can achieve decent performance without actually leveraging the input data. To control for this, we also train a byte-level language model baseline on the same sequence-to-sequence data, excluding the input tokens. If the decompression model is actually learning to decompress Arithmetic Coding, we would expect stronger performance than the byte-level baseline. As we see in Table 5, the baseline model, which does not see the input tokens, has the same performance as the decompression model. [^14] The slight gain is statistically insignificant $(p=0.07)$. Clearly, the models trained for decompression are not actually learning to do decompression. In contrast, we see that models trained on EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}256]$ do appear to learn to perform decompression. An M2 model trained to decompression EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}256]$ compressed data quickly jumps to $90\%$ accuracy, in $9$K updates, and continues to improve to $95\%$ over training. The model is still improving as training ends with no clear signs of overfitting, suggesting the expected $100\%$ accuracy should be achievable with more data. [^15] This could also be explained by how the end of sequence is handled, see Section D.5 It clearly leverages the input as a byte-level language model trained over the same target tokens only reaches an accuracy of $76$.
@@ 6.1 EqualInfoAC is less stable and less semantic than SentencePiece
First, we observe that SentencePiece produces a stable text $\rightarrow$ token mapping. For example, “elephants” appears three times in the sentence, and maps stably to the same two-token sequence in all cases: [ elephant][s]. Similarly, both occurrences of “African” map to the same token: [ African]. By comparison, the EqualInfoAC tokenization is relatively unstable, with each occurrence of these words being segmented in a different way and yielding a different token sequence.
@@
Despite these differences, there is an important similarity between SentencePiece and EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}65\text{k}]$: they are both stable in the token $\rightarrow$ text direction. That is, a given token ID, e.g., token #$500$, will always map to the same output text. This “transparent decoding” property likely makes it easier for a downstream model to learn over these tokens. [^18] Padding to reach a specific window size can require extra computation to discern between padding and characters that compress to all zeros, however we find in Section D.5 that it is not an issue for M2 models.
@@
When we move to versions of EqualInfoAC that contain multiple tokens per window, such as EqualInfoAC$[b\mathord{=}16,\,v\mathord{=}256]$, this transparency is destroyed for all non-initial tokens within a window. This is illustrated in Table 8. When the same token appears window-initially in different contexts, we see the window text has a stable prefix—e.g., token #$151$ always maps to the prefix “le-”. However, when occurring as the second token within a two-token window, there are no apparent correspondences between window text. [^19] A repeated text substring that happens to be aligned with a window multiple times is one of the few cases where the second token will represent the same text. As EqualInfoAC window length increases, the proportion of tokens that are stable decreases. This may explain the observed difficulty of learning over longer windows. The window text for all instances of these tokens can be seen in Section D.8.
@@
Note that Table 8 examines window $\rightarrow$ text, as opposed to token $\rightarrow$ text correspondences. This is because for multi-token windows, the mapping from tokens to text is not well defined. More specifically, each character maps to a particular subsequence of the compressed bitstream, but these may not align with token boundaries. [^20] This can be a source of instability, even in window-initial tokens, see Appendix C. Fig. 7 illustrates the mapping between characters, bits, and tokens. We find that many windows contain a character (shown in purple) whose bits are split across two $8$-bit tokens.
Fig. 7 also highlights that window-initial characters are not being well compressed, with the window-initial token often only covering one or two characters. This is due to our EqualInfoAC procedure fully resetting M1’s context at every window boundary. With no context, M1 cannot make confident predictions, leading to more bits being needed to represent the initial character. While this hurts the overall compression ratio, there is a benefit for learnability in that each window can be decoded in isolation. In Section A.2, we experiment with maintaining M1 context across two or more windows, and find that the improvement in compression ratio does not justify the degradation in bits/byte.
@@ 6.2 AC decoding is learned step-by-step
The second trend concerns the accuracy reached at each position. Here, we observe an increase in accuracy from #1 < #2 < #3, followed by a decrease from #3 < #4 < #5 and so on. [^22] The final token #8 also fits this trend when looking at the increase over non-trivial accuracy. The raw accuracy is higher than previous tokens, #4–7, due to the skewed distribution induced by window-final padding. We interpret the increase across the first three positions as due to the benefit of extra leftward context. [^23] While suggests that $24$ bit windows could be effective, but in Section A.1 we found it is not. This is akin to the initial byte in a word being harder to predict than the following bytes. The decreasing performance at tokens #4 and beyond suggests the model is unable to track AC decompression indefinitely. While the model clearly learns to decompress longer sequences as training progresses, reliably decoding past $32$ bits of AC output appears to be a challenge.
@@ 6.3 Learnable distributions are less uniform
A well-known result in the compression literature is that there can be no recursive compression (Mahoney, 2013). The compression algorithm removes information captured by its model, resulting in a uniform output that appears random to the original model. However, our setting is not recursive compression. Instead, a separate and larger model is trained on the compressed output, which should be able to capture new patterns in the bitstream.
Despite this, the output of compression using M1 appears very uniform, as evidenced by the minimal gains from modeling the unigram token distribution in Table 3. Therefore, it seems reasonable that this uniformity could make it hard for M2 to learn (as all patterns must be contextual). We investigate this by plotting the KL divergence (Kullback & Leibler, 1951) between the observed empirical distribution and a uniform distribution for different segmentations of the bitstream. If the underlying distribution of bits was truly random and independent, then the distribution of unigrams for some bitstream segmentation should remain uniform as $p(b_{i},\dots,b_{i+n})=\prod_{j=i}^{i+n}(p(b_{j}))$ and therefore the KL divergence should remain close to zero. On the other hand, if the distribution diverges from uniform, there is contextual information to be learned when training an LLM to model $p(b_{n}|b_{i},\dots,b_{i+n-1})$.
@@
As a baseline, we used the cryptographic secrets package in Python to generate bitstreams that should be truly random and independent. As such, the KL divergence should remain at 00 when segmented in the same way as the compressed data. The reason this does not hold in Fig. 9 is that the maximum likelihood estimate of entropy, $\hat{H}=-\sum_{x\in\hat{\mathcal{X}}}\hat{p}(x)\log_{2}\hat{p}(x)$, is negatively biased (Paninski, 2003). In Fig. 15 we see that when using a Miller-Madow estimator (Miller, 1955) to correct for this bias, the expected KL of 00 is well within sampling noise bounds. To account for noise in the entropy estimation, we plot $90$th percentile intervals of the KL divergence between the observed entropy from $100$ disjoint samples of the data and the uniform distribution. [^24] As the number of bits in a segmentation grow, the vocabulary size increases exponentially, requiring many more samples. Thus we expect noise in the entropy estimate to grow with $n$. This holds, but it is obfuscated by the log scaling in Fig. 9. In fact, the magnitude of the noise for settings such as GZip and EqualInfoAC is larger than for AC or RNG. This noise behavior is seen in Fig. 14. See Section D.6 for more information on entropy estimation and bias correction.
@@ 7 Conclusion
We have shown there is promise in the idea of training LLMs over neurally compressed text. In the best case, this will allow training over text that is better compressed than standard subword token sequences, while maintaining learnability. This an appealing prospect, as models that read and write more text per token are more efficient to train and serve, and can model longer dependencies.
@@
While bespoke compression methods have developed around different modalities (e.g., text, audio, images, video) and different applications (e.g., delta-of-delta for regular repeating timestamps (Pelkonen et al., 2015)), to our knowledge, no efficient compression methods have been designed specifically for use as LLM tokenizers. We are optimistic that future work will create such methods. Compared to today’s subword tokenizers, we expect these methods (i) will deliver higher compression rates, (ii) will come closer to equal information per token, thus allocating compute more effectively, and (iii) will give models a more direct view of the underlying raw text, thus helping on spelling and pronunciation tasks. As a tradeoff, we expect these neural tokenizers will be somewhat less stable in their text $\leftrightarrow$ token mapping, but perhaps not so unstable as our approach here. In particular, we think it is worth exploring methods under which a given word typically maps to a relatively small number (tens not thousands) of relatable token sequences.